You toss a coin straight upward in a train which is moving at a constant velocity. If the air friction is negligible, the coin will land O (a) In front of you O b) behind you O (c) on your hand O d) at either in front of you or behind you (b) The position of an ol bject is given here t is measured in seconds and X is in meters. Find the average velocity of the object in the time interval t 2s to t 3s. a) 1m/s O b) 2m/s (c) 3m/s (d) 4m/s
Answer #1

as coin is tossed upward but you will be moved forward.

hence coin will land behind you.

(B) x = t^2 - 3t + 2

x(2) = 2^2 - 3(2) + 2 = 0

x(3) = 2 m/s

<v> = x(3) - x(2) / (3 - 2)

= 2 m/s